Jane purchased 100 shares of Zodcom on March 1, 2011 for $131.52 per share. The company...

  1. Jane purchased 100 shares of Zodcom on March 1, 2011 for $131.52 per share. The company instituted a 3:1 stock split 3 months later, and Jane then sold her shares for $49.03 per share. What was Jane's holding period return on this investment?

    56.08%

    10.59%

    11.84%

    37.28%

1 points   

QUESTION 2

  1. Jane purchased 100 shares of Zodcom on March 1, 2011 for $123.81 per share. The company instituted a 3:1 stock split 7 months later, and Jane then sold her shares for $46.33 per share. What was Jane's holding period return on this investment?

    37.42%

    23.89%

    10.92%

    12.26%

1 points   

QUESTION 3

  1. Magdalena purchased 100 shares of BQM on 1/1/2014 for $158.59 per share. During the year, the stock paid a dividend of $0.96 on 2/15/2014, another dividend of $0.96 on 5/15/2014, another dividend of $1.02 on 8/15/2014, and another dividend of $1.02 on 11/15/2014. Magdalena sold the shares on 12/31/2014 for $166.52 per share. What was Magdalena's holding period return for this investment?

    5.64%

    7.14%

    7.50%

    4.80%

1 points   

QUESTION 4

  1. Magdalena purchased 100 shares of BQM on 1/1/2014 for $117.29 per share. During the year, the stock paid a dividend of $0.71 on 2/15/2014, another dividend of $0.71 on 5/15/2014, another dividend of $0.76 on 8/15/2014, and another dividend of $0.76 on 11/15/2014. Magdalena sold the shares on 12/31/2014 for $145.44 per share. What was Magdalena's holding period return for this investment?

    24.64%

    21.37%

    26.50%

    25.20%
Answer #1

Question 1

Investor's total return comprises of 2 elements-capital gain/loss (change in market price) and dividend payment.

HPR = (Sale value - cost of acquisituion + coupon pay) / cost of acquisituion

= ((3*100*49.03)-(131.52*100))/(131.52*100)*100

= (14709-13152)/13152 *100

= 1557/13152 *100

= 11.84%

Question 2

HPR = (Sale value - cost of acquisituion + coupon pay) / cost of acquisituion

= ((3*100*46.33)-(123.81*100))/(123.81*100)*100

= (13899-12381)/12381*100

= 1518/12381*100

= 12.26%

Note: As per HOMEWORKLIB RULES, when multiple questions are posted, we are mandated do only the first question. So please post the questions individually.